# Importance Ranking in Complex Networks via Influence-aware Causal Node Embedding

### v2

@@ 2.2 Causal representation learning
Causal relationships have the ability to capture the fundamental mechanism of data generation and are stable across various contexts [jiang2025integrating, zhou2025vulnerability]. Learning causal representations is significant for enhancing the robustness of predictive models. Zheng et al. [zheng2018dags] introduced NOTEARS, a novel method that reformulates the DAG discovery problem as a continuous optimization task incorporating acyclicity constraints, enabling solution via numerical techniques. Subsequently, Ng et al. [ng2019graph] expanded upon NOTEARS within a graph autoencoder framework, leveraging multilayer perceptrons (MLPs) to capture and model nonlinear structural equations, thereby advancing the capability to handle complex, non-linear relationships in causal inference. Yang et al. [yang2021learning] integrated a deep autoencoder and a causal structure learning model to learn causal representations using data. Yu et al. [yu2019dag] proposed DAG-GNN, a deep generative model that uses a graph neural network-based variational autoencoder to learn DAG structures from data, effectively generalizing linear structural equation models to capture nonlinear relationships and handle diverse variable types.
Most existing causal representation learning methods aim to uncover causal relationships among variables, but they are not tailored for downstream ranking tasks, making it difficult to ensure the learned representations are effective for ranking. To this end, ICAN integrates a causal ranking loss into its representation learning model, enabling it to produce robust, low-dimensional node embeddings well suited for node-importance ranking tasks.
@@ 3.1 Problem statement
Given a graph $G=\{V,E\}$, where $V$ is the set of nodes and $E$ is the set of edges. The notation $v_{i}\in V$ denotes a node in $V$, and $e_{ij}\in E$ denotes an edge from node $v_{i}$ to $v_{j}$. The graph can be represented by an $n\times n$ adjacency matrix $\bm{A}$, where $n$ denotes the number of nodes. The entries of the matrix are defined such that $\bm{A}_{ij}=1$ when $e_{ij}\in E$, and $\bm{A}_{ij}=0$ otherwise. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix. For graphs without node attributes, $\bm{X}$ can be obtained by applying a node embedding method (e.g., Node2Vec [grover2016node2vec]) to the adjacency matrix $\bm{A}$.
The node importance ranking is a critical problem in network analysis. Given the graph $G=\{V,E\}$, our task is to assign a score $s_{i}$ to each node $v_{i}\in V$, forming a score vector $\bm{S}=(s_{1},s_{2},\dots,s_{n})$, These scores are then used to derive a node importance ranking $\bm{R}=(r_{1},r_{2},\dots,r_{n})$, where $r_{i}$ denotes the rank of node $v_{i}$. The ranking $\bm{R}$ satisfies the following: $r_{i}=r_{j}$ if $s_{i}=s_{j}$, and $r_{i}<r_{j}$ if $s_{i}<s_{j}$, meaning that a higher score corresponds to a higher (i.e., numerically lower) rank. Unlike methods that emphasize the absolute magnitude of scores, our approach focuses on preserving the relative ordering of nodes consistent with their underlying importance, making ranking consistency the primary objective.
@@ 3.2 Assumptions
(Causal graph [pearl2014probabilistic]) For Bayesian Network $U,\mathbb{G},P>$, where $U$ denotes the set of variables, $\mathbb{G}$ is a Directed Acyclic Graph (DAG) on $U$ and $P$ denotes the probability distribution of $U$, then Bayesian Network $U,\mathbb{G},P>$ can be used to express the causal relationships between the variables in $U$. In DAG $\mathbb{G}$, for a pair of directly connected parent-child variables, the parent variable is the direct cause of the child variable, and the child variable is the direct outcome of the parent variable. We assume that the causality between variables can be expressed by $\mathbb{G}$ as a causal graph.
@@
(Causal Markov [fields1990independence]) A variable $X$ is independent of every other variable (except X’s effects) conditional on all of its direct causes.
@@
(Faithfulness [guo2023adaptive]) Every conditional independence that holds in the distribution $P$ is entailed by the Markov condition applied to $\mathbb{G}$. Faithfulness ensures that all conditional independencies in the data correspond to missing edges in the causal graph.
@@ 4.1 Overview of the method
The framework of ICAN is illustrated in Fig. 1. ICAN consists of two core modules: the causal node embedding module and the causal ranking prediction module. The causal representation learning module aims to learn low-dimensional node embeddings that are causally related to node importance. Under Assumption 1, the causality among node embeddings can be described by a DAG $\mathbb{G}$. Upon obtaining the $p$-dimensional $(p\leq d)$ feature representation, the node influence score variable $\bm{Y}$, which can be used to characterize the node importance, is introduced as an additional node to form a new hidden layer. The adjacency matrix $\bm{W}=\{w_{ij}\}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$, which encodes the causal relationships among the low-dimensional node embeddings $\bm{H}^{(l)}$, is then integrated into the autoencoder framework to enable message passing under causal mechanisms. By optimizing $\bm{W}$, ICAN can learn optimal representations that are causally linked to node importance. The learned latent representations are then fed into the causal ranking prediction module, where the Markov Blanket (MB) of the node importance score variable is leveraged to predict the ranking. In this way, ICAN can produce robust, network-invariant node embeddings, enabling cross-network importance ranking predictions. In the following, we provide a detailed description of the ICAN framework.
Figure 1: The framework outline of ICAN.
@@ 4.2 Causal node embedding module
Here, $\widetilde{\bm{A}}=\bm{D}^{-\frac{1}{2}}(\bm{A}+\bm{I})\bm{D}^{-\frac{1}{2}}$ and $\bm{D}$ is the degree matrix. $l$ denotes the number of hidden layers. M[:,1:(p1)]M[:,1:({p}-1)] denotes the matrix composed of the first p1p-1 columns. Let $\bm{\varPhi}^{(l)}=\hat{\bm{A}}$ be the reconstructed adjacency matrix. $\bm{w}_{1}^{(i)}$ and $\bm{b}_{1}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ encoding layer, respectively. Similarly, $\bm{w}_{2}^{(i)}$ and $\bm{b}_{2}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ decoding layer, respectively. $\bm{H}^{(m-1)}\in\mathbb{R}^{n\times p}$ in the encoding process can be seen as the low-dimensional representations.
@@
where the adjacent matrix $\bm{W}$ is required to satisfy the constraint in Eq. (18). $g_{1}$ and $g_{2}$ can be regarded as two nonlinear functions used to describe the encoding and decoding process, respectively. Therefore, it is noted here that the causal node embedding module can be seen as a graph autoencoder.
Based on the causal graph $\mathbb{G}$ or its corresponding adjacency matrix $\bm{W}$, the Markov Blanket $\text{MB}(\bm{Y})$ of the node influence score variable $\bm{Y}$ can be identified under the causal Markov assumption (Assumption 2) and the faithfulness assumption (Assumption 3). This MB($\bm{Y}$) constitutes the set encompassing the parent, child, and spouse nodes of the node influence score variable $\bm{Y}$. The robust causal relationships between the node influence score variable $\bm{Y}$ and the features within its Markov Blanket, $\text{MB}(\bm{Y})$—designated as the causal feature subset—reflect an invariant influence propagation mechanism across networks. As such, utilizing this invariant subset for node-level prediction tasks enhances model performance and generalization capability on unseen networks.
@@ 4.3 Causal ranking prediction module
Here, $\bm{A}$ represents the adjacency matrix of the graph, and $\bm{w}_{3}^{(i)}$ and $\bm{b}_{3}^{(i)}$ are the weight matrix and bias matrix respectively.
Task-aware ranking loss function. We introduce a new ranking loss function, CausalListMLE, which can be viewed as an extension of ListMLE [xia2008listwise] for optimizing the overall ranking prediction module. This list-wise ranking loss directly maximizes the likelihood of the entire ranked permutation, rather than focusing on individual scores as in point-wise losses like MSE. By modeling the ranking distribution, CausalListMLE focuses on preserving the correct relative order among all items, and typically achieves better performance in ranking-based learning tasks. The causal ranking loss $\mathcal{L}_{2}$ is defined as
@@
where $x_{\pi(i)}$ is the predicted score of the node at position $i$ in the ranking $\pi$. By maximizing this likelihood, the causal ranking loss encourages the model to produce ranking that closely align with the influence score.
@@ 4.4 The objective function
We have designed the causal reconstruction loss $\mathcal{L}_{1}$ to ensure the learned representations accurately causal dependencies underlying node importance, and the causal ranking loss $\mathcal{L}_{2}$ to optimize them for the downstream ranking task. In addition, to mitigate model overfitting and prevent performance degradation, we design a regularization term $\mathcal{L}_{3}$ on the weight parameters to the objective function $\mathcal{L}$:
@@
To ensure that the learned $\bm{W}$ forms a DAG, this objective must satisfy the acyclicity constraint [zheng2018dags]:
@@
The adjacency matrix $W$ corresponds to an acyclic graph $\mathbb{G}$ if and only if Eq. (18) holds.
@@
To solve above equality-constrained problem, the augmented Lagrangian method [ng2019graph] is employed to convert it into an unconstrained problem. After introducing the Lagrange multiplier and the penalty term, the objective function Eq. (4.4) is reformulated as the following augmented Lagrangian function:
@@
where $\alpha$ denotes the Lagrange multiplier, $\rho>0$ denotes the penalty parameter and $h(\bm{W})=\text{tr}(e^{\bm{W}\odot\bm{W}})-(p+1)$. We then have the following update rules for the adjacent matrix $\bm{W}$, $\alpha$ and $\rho$:
@@
$\rho^{(t+1)}$ $=\begin{cases}\beta\rho^{(t)},&\mbox{if }|\bm{W}^{(t+1)})|\leq\theta|h(\bm{W}^{(t)})|,\\ \rho^{(t)},&\text{otherwise},\end{cases}$ (24)
@@
4:  Update the adjacent matrix $\bm{W}^{(t+1)}$ by Eq. (22);

### v3

@@ 2.3 Causal representation learning
Causal relationships have the ability to capture the fundamental mechanism of data generation and are stable across various contexts, mitigating the impact of confounding factors [31, 84]. Learning causal representations is significant for enhancing the robustness of predictive models. Zheng et al. [80] introduced NOTEARS, a novel method that reformulates the DAG discovery problem as a continuous optimization task incorporating acyclicity constraints, enabling solution via numerical techniques. Subsequently, Ng et al. [51] expanded upon NOTEARS within a graph autoencoder framework, leveraging multilayer perceptrons (MLPs) to capture and model nonlinear structural equations, thereby advancing the capability to handle complex, non-linear relationships in causal inference. Yang et al. [71] integrated a deep autoencoder and a causal structure learning model to learn causal representations using data. Yu et al. [73] proposed DAG-GNN, a deep generative model that uses a graph neural network-based variational autoencoder to learn DAG structures from data, effectively generalizing linear structural equation models to capture nonlinear relationships and handle diverse variable types. Wu et al. [66] proposed DIR that constructs intrinsically interpretable graph neural networks by conducting interventions on the training distribution to create interventional distributions, thereby identifying causal rationales invariant across distributions while filtering out spurious patterns for improved out-of-distribution generalization and interpretability. Fan et al. [15] proposed DisC that learns disentangled causal and bias substructures by employing a parameterized edge mask generator to split input graphs into causal and bias subgraphs, training separate GNN modules with causal-aware and bias-aware loss functions, and generating counterfactual unbiased samples to decorrelate the variables for improved generalization. Chen et al. [8] proposed CIGA that extracts invariant subgraphs containing causal information about labels to learn causally invariant representations for guaranteed out-of-distribution generalization on graphs under various distribution shifts. Mo et al. [48] proposed GCIL, a causal-inspired graph contrastive learning method that uses spectral graph augmentation to intervene on non-causal factors and incorporates invariance and independence objectives to extract causal information for invariant representation learning. Most of these causal representation learning methods aim to uncover causal relationships among variables, but they are not tailored for downstream ranking tasks, making it difficult to ensure the learned representations are effective for ranking.
We summarize the key lessons learned from the related work as follows:
  • Most existing node importance ranking methods rely heavily on network topology and decouple representation learning from ranking objectives, which limits their generalization across different networks.
  • Task-oriented graph representation learning improves feature–task alignment, but remains largely correlation-driven and vulnerable to spurious, network-specific patterns.
  • Causal representation learning offers invariant and robust representations, yet has rarely been explored in the context of node importance ranking.
Motivated by these observations, ICAN integrates a causal ranking loss into the representation learning process, enabling the model to learn robust, low-dimensional node embeddings that are generalizable across diverse networks and well suited for node importance ranking tasks.
@@ 3.1 Problem statement
Given a graph $G=\{V,E\}$, where $V$ is the set of nodes and $E$ is the set of edges, the notation $v_{i}\in V$ denotes a node in $V$, and $e_{ij}\in E$ denotes an edge from node $v_{i}$ to $v_{j}$. The graph can be represented by an $n\times n$ adjacency matrix $\bm{A}$, where $n$ denotes the number of nodes. The entries of the matrix are defined such that $\bm{A}_{ij}=1$ if $e_{ij}\in E$, and $\bm{A}_{ij}=0$ otherwise. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denote the feature matrix. For graphs without node attributes, $\bm{X}$ can be obtained by applying a node embedding method (e.g., node2vec [22]) to the adjacency matrix $\bm{A}$.
The node importance ranking is a critical problem in network analysis. Given the graph $G=\{V,E\}$, our task is to assign a score $s_{i}$ to each node $v_{i}\in V$, forming a score vector $\bm{S}=(s_{1},s_{2},\dots,s_{n})$, These scores are then used to derive a node importance ranking $\bm{R}=(r_{1},r_{2},\dots,r_{n})$, where $r_{i}$ denotes the rank of node $v_{i}$. The ranking $\bm{R}$ satisfies the following: $r_{i}=r_{j}$ if $s_{i}=s_{j}$, and $r_{i}<r_{j}$ if $s_{i}>s_{j}$, meaning that a higher score corresponds to a higher (i.e., numerically lower) rank. Unlike methods that emphasize the absolute magnitude of scores, our approach focuses on preserving the relative ordering of nodes consistent with their underlying importance, making ranking consistency the primary objective.
@@ 3.2 Assumptions
(Markov condition [16]) Given a Directed Acyclic Graph (DAG) $\mathbb{G}$, each variable $X$ in $\mathbb{G}$ is independent of any subset of its non-descendants conditioning on its parents.
Definition 1.
The triplet $U,\mathbb{G},P>$, where $U$ denotes the set of variables, $\mathbb{G}$ is a DAG on $U$, and $P$ denotes the probability distribution of $U$, is called a Bayesian network iff it satisfies the Markov condition.
@@
(Causal graph [56]) For Bayesian Network $U,\mathbb{G},P>$, it can be used to express the causal relationships between the variables in $U$. In DAG $\mathbb{G}$, for a pair of directly connected parent-child variables, the parent variable is the direct cause of the child variable, and the child variable is the direct outcome of the parent variable. We assume that the causality between variables can be expressed by $\mathbb{G}$ as a causal graph.
Definition 2.
($d$-separation [36, 55]) In a DAG $\mathbb{G}$, a path $\pi$ is said to be blocked by a set of nodes $C$ if and only if (i) $\pi$ contains a chain $V_{i}\rightarrow V_{k}\rightarrow V_{j}$ or a fork $V_{i}\leftarrow V_{k}\rightarrow V_{j}$ such that the middle node $V_{k}$ is in $C$, or (ii) $\pi$ contains a collider $V_{i}\rightarrow V_{k}\leftarrow V_{j}$ such that the middle node $V_{k}$ is not in $C$ and such that no descendant of $V_{k}$ is in $C$. Furthermore, in a DAG $\mathbb{G}$, we say that two disjoint subsets of vertices $A$ and $B$ are d-separated by a third (also disjoint) subset $C$ if every path between nodes in $A$ and $B$ is blocked by $C$. We then write $A\perp_{\mathbb{G}}B\,|\,C.$
@@
(Faithfulness [55, 24]) Consider a distribution $P$ and a DAG $\mathbb{G}$. $P$ is faithful to the DAG $\mathbb{G}$ if
$A\perp B\,|\,C\implies A\perp_{\mathbb{G}}B\,|\,C$
for all disjoint vertex sets $A,B,C$.
Definition 3.
(Markov Blanket (MB)) Consider a DAG $\mathbb{G}$ and a target node $Y$. Then, the Markov Blanket MB of $Y$ includes its parents, its children, and the parents of its children.
Faithfulness ensures that all conditional independencies in the data correspond to $d$-separations in the causal graph. Under this assumption and Markov condition, all other features that are not in the Markov blanket set of the target variable (consisting of the parents, children, and spouses) are conditionally independent of the target variable, conditioning on the MB set. This means that the information of the non-Markov blanket feature representations is blocked given the Markov blanket feature representations [71].
@@ 4.1 Overview of the method
The framework of ICAN is illustrated in Fig. 1. ICAN consists of two core modules: the causal node embedding module and the causal ranking prediction module. The causal node embedding module aims to learn low-dimensional node embeddings that are causally related to node importance. Under Assumption 1, the causality among node embeddings can be described by a DAG $\mathbb{G}$. Upon obtaining the $p$-dimensional $(p\leq d)$ feature representation, the node influence score variable $\bm{Y}$, which can be used to characterize the node importance, is introduced as an additional node to form a new hidden layer. The adjacency matrix $\bm{W}=\{w_{ij}\}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$, which encodes the causal relationships among the low-dimensional node embeddings $\bm{H}^{(l)}$, is then integrated into the autoencoder framework to enable message passing under causal mechanisms. By optimizing $\bm{W}$, ICAN can learn optimal representations that are causally linked to node importance. The learned latent representations are then fed into the causal ranking prediction module, where the Markov blanket of the node importance score variable is leveraged to predict the ranking. In this way, ICAN can produce robust and network-invariant node embeddings, enabling cross-network importance ranking predictions. In the following, we provide a detailed description of the ICAN framework.
Figure 1: The framework outline of ICAN. The model takes the adjacency matrix $\bm{A}$ as inputs and processes them through two key components: (1) Causal node embedding module, which learns low-dimensional node embeddings that are causally related to node importance; and (2) Causal ranking prediction module, which utilizes the Markov Blanket of the node influence score variable to predict the final ranking list. This model, trained solely on synthetic networks, generalizes effectively to diverse real-world graphs for node importance ranking.
@@ 4.2 Causal node embedding module
Here, $\widetilde{\bm{A}}=\bm{D}^{-\frac{1}{2}}(\bm{A}+\bm{I})\bm{D}^{-\frac{1}{2}}$ and $\bm{D}$ is the degree matrix. $l$ denotes the number of hidden layers. M[:,1:(p1)]M[:,1:({p}-1)] denotes the matrix composed of the first p1p-1 columns. Let $\bm{\varPhi}^{(l)}=\hat{\bm{A}}$ be the reconstructed adjacency matrix. $\bm{w}_{1}^{(i)}$ and $\bm{b}_{1}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ encoding layer, respectively. Similarly, $\bm{w}_{2}^{(i)}$ and $\bm{b}_{2}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ decoding layer, respectively. The dimensions of the intermediate matrices are defined as $\bm{H}^{(i)}\in\mathbb{R}^{n\times p}$ and $\bm{\varPhi}^{(i)}\in\mathbb{R}^{n\times(p+1)}$ for $i\in\{0,\dots,m-1\}$; $\bm{H}^{(i)}\in\mathbb{R}^{n\times(p+1)}$ for $i\in\{m,\dots,l\}$; and $\bm{\varPhi}^{(m)}\in\mathbb{R}^{n\times p}$, $\bm{\varPhi}^{(l)}\in\mathbb{R}^{n\times n}$. Within this structure, $\bm{H}^{(m-1)}\in\mathbb{R}^{n\times p}$ in the encoding process can be seen as the low-dimensional representations.
@@
where the adjacency matrix $\bm{W}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$ is required to satisfy the constraint in Eq. (18). $g_{1}$ and $g_{2}$ can be regarded as two nonlinear functions used to describe the encoding and decoding process, respectively. Therefore, it is noted here that the causal node embedding module can be seen as a graph autoencoder.
Utilizing the learned adjacency matrix $\bm{W}$ associated with the DAG $\mathbb{G}$, the Markov blanket of the node influence variable $\bm{Y}$ can be precisely identified. Under the Markov condition (Assumption 1) and Faithfulness assumptions (Assumption 3), $Y$ is conditionally independent of all other variables in $\mathbb{G}$ given its Markov blanket. Generally, $\text{MB}(\bm{Y})$ constitutes the minimal sufficient set that shields $\bm{Y}$ from all other non-descendant nodes. In practice, while the direct causal dependencies are manifested as non-zero entries in the row and column of $\bm{W}$ associated with $\bm{Y}$, the complete MB set is extracted by traversing the topological connectivity of $\mathbb{G}$. By isolating these causal feature indices, we effectively filter out task-irrelevant embeddings that exhibit only spurious correlations with the target. Consequently, the low-dimensional node embeddings in $\bm{H}^{(m)}$ can be divided into two groups: causal node embeddings $\text{MB}(\bm{Y})$, and the other feature representations which may be task-irrelevant features. As such, utilizing this invariant subset for node-level prediction tasks enhances model performance and generalization capability on unseen networks.
@@ 4.3 Causal ranking prediction module
Here, $\bm{A}$ represents the adjacency matrix of the graph, and $\bm{w}_{3}^{(i)}$ and $\bm{b}_{3}^{(i)}$ are the weight matrix and bias matrix respectively, $\bm{E}^{(i)}\in\mathbb{R}^{n\times p}$.
Task-aware ranking loss function. We introduce a new ranking loss function, CausalListMLE, which can be viewed as an extension of ListMLE [67] for optimizing the overall ranking prediction module. This list-wise ranking loss directly maximizes the likelihood of the entire ranked permutation, rather than focusing on individual scores as in point-wise losses like MSE. By modeling the ranking distribution, CausalListMLE focuses on preserving the correct relative order among all items, and typically achieves better performance in ranking-based learning tasks. The causal ranking loss $\mathcal{L}_{2}$ is defined as
@@
where $x_{\pi(i)}$ is the predicted score of the node at position $i$ in the ranking $\pi$. By maximizing this likelihood, the causal ranking loss encourages the model to produce a ranking that closely aligns with the influence score.
Remark: It is essential to clarify that while the mathematical formulation of CausalListMLE (Eq. (14)) shares the same Plackett–Luce likelihood structure as standard ListMLE [67], our CausalListMLE derives its novelty from a fundamental difference in the input and the resulting optimization objective. Standard ListMLE computes ranking scores $f(\bm{X})$ using all input features, often leading to models that overfit spurious correlations. In contrast, CausalListMLE is fundamentally designed to restrict the input of the scoring function to the causal feature subset $\text{MB}(\bm{Y})$ (the Markov Blanket of the node influence). This restriction enables the model to co-optimize the alignment between the causal embeddings and the ranking order, thereby yielding invariant and importance-related features, which are critical for achieving robust cross-network generalization.
@@ 4.4 The objective function
We have designed the causal reconstruction loss $\mathcal{L}_{1}$ to ensure the learned representations accurately capture the causal dependencies underlying node importance, and the causal ranking loss $\mathcal{L}_{2}$ to optimize them for the downstream ranking task. In addition, to mitigate model overfitting and prevent performance degradation, we design a regularization term $\mathcal{L}_{3}$ on the weight parameters to the objective function $\mathcal{L}$:
@@
Remark: The term $\mathcal{L}_{3}$ corresponds to the standard $L^{2}$ (weight-decay) regularization applied to model weights. This penalty constrains the magnitude of the weight parameters by adding their squared $\ell_{2}$ norm to the loss, which effectively reduces overfitting by preventing excessively large weights. Since each bias parameter controls only a single variable, not regularizing it does not lead to significant variance. Furthermore, regularizing bias parameters can result in noticeable underfitting. Therefore, in neural networks, regularization is typically applied only to weight parameters, not bias parameters [21]. Also, in this paper, the regularization term $\mathcal{L}_{3}$ does not impose constraints on bias terms.
@@
To ensure that the learned $\bm{W}$ forms a DAG, this objective must satisfy the acyclicity constraint [80]:
@@
The adjacency matrix $\color[rgb]{0,0,0}\bm{W}$} corresponds to an acyclic graph $\mathbb{G}$ if and only if Eq. (18) holds.
@@
To solve above equality-constrained problem, the augmented Lagrangian method [51] is employed to convert it into an unconstrained problem. After introducing the Lagrange multiplier and the penalty term, the objective function in Eq. (4.4) is reformulated as the following augmented Lagrangian function:
@@
where $\alpha$ denotes the Lagrange multiplier, $\rho>0$ denotes the penalty parameter and $h(\bm{W})=\text{tr}(e^{\bm{W}\odot\bm{W}})-(p+1)$. We then have the following update rules for the adjacency matrix $\bm{W}$, $\alpha$ and $\rho$:
@@
$\rho^{(t+1)}$ $=\begin{cases}\beta\rho^{(t)},&\mbox{if }|h(\bm{W}^{(t+1)})|\leq\theta|h(\bm{W}^{(t)})|,\\ \rho^{(t)},&\text{otherwise},\end{cases}$ (24)
@@
4:  Update the adjacency matrix $\bm{W}^{(t+1)}$ by Eq. (22);